OpenStudy (anonymous):

3sqrt{y^13} 3sqrt{81y^14} 3sqrt{y^13} 3sqrt{81y^14} @Mathematics

OpenStudy (anonymous):

\[3\sqrt{y ^{13}} 3\sqrt{81y ^{14}}\]

OpenStudy (anonymous):

\[81\sqrt{y^{27}}\]

OpenStudy (anonymous):

\[3y ^{9}*\sqrt[3]{3}\]

OpenStudy (anonymous):

im not getting those. how did you get the answer

OpenStudy (anonymous):

\[3\sqrt{y^{13}3\sqrt{81y^{14}}}\]

OpenStudy (anonymous):

Which is the correct question?

OpenStudy (anonymous):

\[\sqrt[3]{y^{13}\sqrt[3]{81}}\]

OpenStudy (anonymous):

the one that i typed in as the equation at the very top

OpenStudy (anonymous):

i have to multiply the two

OpenStudy (anonymous):

\[\sqrt[3]{y ^{13}}*\sqrt[3]{81y ^{14}}\] \[y ^{13/3}*y ^{14/3}*81^{1/3}\] \[y ^{27/3}*\sqrt[3]{27}*\sqrt[3]{3}\] \[y ^{9}*3*\sqrt[3]{3}\]

OpenStudy (anonymous):

\[y ^{9}*3^{4/3}\]

OpenStudy (anonymous):

\[81 y^{27/2} \]Replace y by 0.7 in the following list and evaluate each list element:\[\left\{81 \sqrt{y^{13}} \sqrt{y^{14}},81 y^{27/2}\right\}\text{ /. } y\to 0.7 \]\[\{0.656612,0.656612\} \]

OpenStudy (anonymous):

Y is a variable, it can have any value

OpenStudy (anonymous):

Not a complete "proof", but I claim that the second list element is a valid simplification of the problem expression.